OpenStudy (anonymous):

how many gigabytes are in a terabyte recall that a gigabyte is 2^30 bytes and a terabyte is 2^40 bytes. Divide the powers to find the answer.

hartnn (hartnn):

so what u get by \(\large \dfrac{2^{40}}{2^{30}}=...?\)

hartnn (hartnn):

use, \(\huge \dfrac{x^m}{x^n}=x^{m-n}\)

OpenStudy (anonymous):

Well since that bases are the same couldnt i just divide the exponents?

OpenStudy (goldphenoix):

You forgot the law of exponent. When you are dividing 2 numbers with the same base, then the base stays the same, but the exponent would be subtracted from the numerator and the denominator. So you can't divide the exponents. Since subtracting and dividing are different.

OpenStudy (anonymous):

So you would subtract them?

OpenStudy (goldphenoix):

Yes, like hartnn said.

OpenStudy (anonymous):

So i have 2^30 - 2^40 = 2^10 so there are 2^10 gigabytes in a terabyte?

OpenStudy (goldphenoix):

For example, if we have 10^3 /10^2. Then in expanded form, it would look like: \[\large \large \frac{ 10 \times 10 \times 10 \times 10 }{ 10 \times 10 } = 10^2 = 100\]

OpenStudy (goldphenoix):

10^4 / 10^2, my bad. Then it would look like 10^2.

hartnn (hartnn):

yes, the answer seems correct :) but it is \(\dfrac{2^{40}}{2^{30}}=2^{40-30}=2^{10}\)